OpenStudy (anonymous):

Simplify the quotient 4.5 x 10^3/9 x 10^7. Write your answer in scientific notation. Show your work.?

OpenStudy (anonymous):

Is the question to simplify\[\frac{4.5\times10^3}{9\times10^7}\]?

OpenStudy (triciaal):

@Calcmathlete I have the same interpretation of the problem

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Ok, assuming that was the question, then you can simplify the problem by splitting up the division. \[\frac{4.5 \times 10^3}{9 \times 10^9} \implies\frac{4.5}{9}\times\frac{10^3}{10^7}\] At this point, you can perform the divisions separately. Do the \(\frac{4.5}{9}\) first, and use your exponent rules to simplify the right fraction. Then, just put the \(\times\) sign to indicate multiplication. Can you finish the problem from there?

OpenStudy (triciaal):

hint 9 = 2*4.5

OpenStudy (anonymous):

do i divide 4.5 from 9

OpenStudy (triciaal):

the standard form for scientific notation is; a*10^n where a is between 0 and 10; n is an integer

OpenStudy (anonymous):

Yes. Divide 4.5 and 9. \(4.5\div9 = ?\)

OpenStudy (triciaal):

yes I am picky with math add and subtract from multiply and divide by

OpenStudy (anonymous):

As @triciaal said above, once you get what 4.5/9 is, that will be the 'a' in the form posted.

OpenStudy (anonymous):

0.5 then 1000/10000000

OpenStudy (triciaal):

rule for exponents when you multiply numbers with the same base add the powers when you divide numbers with the same base subtract the powers

OpenStudy (anonymous):

thank you that makes it less confusing

OpenStudy (anonymous):

Ok, that's right so far. The 0.5 is correct. To get it into the scientific notation form, you have: \(0.5\times10^n~~\)The next step would be to solve for what the n is. Do you remember your exponent rules? \[\frac{10^a}{10^b} = 10^{a-b}\]

OpenStudy (triciaal):